if ax+by+c=0 is the equation of the line then x and y intercept respectively are
 

Grade:12th

1 Answers

Arun
5 years ago
a x + b y +c = 0
ax + by = -c
Now we have to make RHS as 1
ax /-c +  by/-c  = 1
x /(-c/a) +  y/(-c/b)  = 1
Hence
X intercept = -c/a
Y intercept = -c/b